OpenStudy (anonymous):

Math help please. How many solutions does the equation have? 1) 2(2x+5)=4(x+3) A. One solution B. Infinite solutions C. No solution 2) a+5=1/5(15+5a) A. One solution B. Infinite solutions C. No solution

OpenStudy (anonymous):

I accidentally pressed post without finishing. Ill be posting more. Thanks to anyone who will help!

OpenStudy (anonymous):

1.c

OpenStudy (anonymous):

2.c

OpenStudy (anonymous):

3. 4x+1=5+2(2x-4)

OpenStudy (anonymous):

4. Which expressions would complete this equation so that it has one solution? Choose exactly two answers that are correct. 4(2x-3)-4x= A. 4x-12 B. 3x-4 C. 3(2x-6) D. 4(x-5)

OpenStudy (anonymous):

5. What is the best way to classify each equation? 1: 8x+24=2(4x+12) 2: 5x+18-x=2(2x+8) 3: 7(3x-2)=20x-13 4: 3x+2(x-10)=5(x-4) Put A, B, Or C on each number. The same letter can be used several times. A: Identity B: Contradiction C: Neither Thats all the questions. Please awnser!

OpenStudy (anonymous):

(I need this awnsered ASAP please! I might give medals/fan)